OpenStudy (anonymous):

Write -(2/3)^5 as a repeated multiplication. Identify the base and the exponent of your power.

OpenStudy (kropot72):

If the expression is rewritten as follows, the solution should be clearer: \[\large -(\frac{2}{3})^{5}=-1\times(\frac{2}{3})^{5}\]

OpenStudy (kropot72):

@Dianegrace Are you there?